# Compositional Reasoning with Transformers, RNNs, and Chain of Thought

### v1

@@ 4.2 Constant depth transformers cannot solve CRQs
[…]
@@
[…]
@@
[…]
We note that arithmetic problems, as presented in Feng et al. (2024), can also be embedded as certain generalized CRQs. This embedding is presented in Section 7 and Appendix F. Thus, both Theorem 4.1 and Theorem 4.3 can be applied to arithmetic problems, if they are presented in a tree structure. It is in general possible to apply a preprocessing procedure using a depth $O(\log(n))$ transformer to turn an arithmetic problem presented in a sequential form, into a tree structure which captures the order of arithmetic operations.
@@ 5 Solution Using Shallow RNNs
In this section we will show that under certain assumptions, RNNs can solve CRQs. The proofs for this section appear in Appendix C. We first define RNNs in the following way:
@@ 5.1 The order of the nodes matters
Our next result shows that there are bad orderings. This means that we can construct a family of CRQs and order the nodes adversarially so that any RNN must have a hidden dimension of size $\Omega(n)$ to solve them successfully.
@@
The proof is given in Section C.1. Note that Theorem 5.2 doesn’t depend on the depth of the RNN. This means that given certain ordering of the nodes, the RNN must encode many of them into the hidden dimension. Also note that it is easy to construct an RNN with a hidden dimension of size $O(n)$ that solves all CRQs. The idea is just to encode all the nodes in the hidden state, which can be done if we have enough memory, and then use a large enough neural network to solve the CRQ. Finding such a neural network is also doable because $2$-layer networks have the universal approximation property (Cybenko, 1989; Leshno et al., 1993). Since the number of possible outputs is finite (because the alphabet is finite), we can approximate the solution up to a small enough accuracy and then threshold over the output.
The proof of Theorem 5.2 (Section C.1) uses a communication complexity argument. Specifically, we use a reduction from the set disjointness problem (see C.1 for a formal definition). We then construct a communication protocol between two parties, where each one of them has knowledge of only half of the inputs. When using the bad ordering of the nodes, this communication protocol forces the first party to encode all of its inputs into the hidden state when passing it to the second party.
We also note that Feng et al. (2024) proved that a specific construction of RNNs that solve arithmetic problems requires a hidden state of size $\Omega(n)$. Our results are stronger in the sense that: (1) Our lower bound is general and applies to any possible construction, so long as the bad order is used, and (2) We next show that the $\Omega(n)$ memory requirement can be greatly reduced just by re-ordering the inputs.
@@ 5.2 Memory-rank sort
In this subsection, we restrict ourselves to CRQs with full binary trees. That is, the degree of each node is 00 or $2$. This simplifies the proofs and presentation. In Appendix D, we explain how to extend the results from this section to non-binary trees. Also, note that every CRQ with a max degree of $k$ for each node can be converted into an equivalent CRQ with a binary tree, and the number of nodes increases by a factor of at most $2$ (see Section D.1). We also note that a node of degree $1$ is unnecessary in the context of CRQs, because a sub-question of degree $1$ has only one possible answer.
@@
Intuitively, the memory rank is the smallest possible stack size needed for a stack machine to solve the CRQ if we are allowed to pick the order of the nodes. A stack machine is an automaton augmented with a stack that it can uses as memory. It processes the inputs sequentially, optionally pushing and popping a finite number of elements from its stack at each step and performing computations with the resulting vectors. We should order the inputs according to a post-ordering depth-first search of the CRQ, meaning that the position of a node in the ordering corresponds to when the depth first search last visits it. This ordering akin to reverse Polish notation, which notates an arithmetic expression like $(2+3)*6$ as the following sequence: 2 3+ 62\,3\,+\,6\,*2 3 + 6 ∗. By ordering the inputs this way, we ensure that whenever the automaton reads a non-leaf node, the top two elements on the stack are the solutions to the CRQs defined by that node’s two children. It can simply pop them off, compute the solution to the current node, and push the result back onto the stack to save it for later.
Even using a depth-first search ordering, the size of the stack used by this machine could grow as large as $n$ in the worst case. However, by ordering the children of each node carefully, we can reduce the stack depth to only $\log n$. To compute the CRQ corresponding to a given node $v$, we (1) solve for one of its children, (2) save this result on the stack, then (3) solve for the second child. If we solve $c_{1}(v)$ first, then the largest stack depth during the first stage is $\textbf{mr}(c_{1}(v))$ and the largest stack depth during the second stage is $\textbf{mr}(c_{2}(v))+1$. The overall largest stack depth is the maximum of these two. Thus, to use as little memory as possible, we should always start with the child whose memory rank is larger. This sorting algorithm is defined formally in Algorithm 1. See Figure 2 for an example of memory-rank sort of a tree. In Theorem 5.4, we will simulate this stack machine using a RNN with hidden dimension of size $\textbf{mr}(T)$.
@@
Figure 2: Memory-rank sorting of a tree. Each node is labeled with its memory rank, which ranges from 0 to 2 in this case. The number below each node is its ordering according to Algorithm 1. Note that we first traverse the right branch of the tree, since the right child of the root has a higher memory-rank than the left child.
@@ 5.3 Shallow RNNs can solve CRQs
For any $n\in\mathbb{N}$ there exists an RNN with $5$ layers and hidden dimension $O(d\log(n))$ that solves any CRQ with a binary tree, if the nodes are ordered by Algorithm 1
The proof is given in Section C.2. The main idea of the proof is to simulate a stack machine with the RNN. The stack will contain the vectors $\mathbf{x}_{i}$ that are needed for future calculations, and will have two possible operations: (1) pop out vectors, and (2) push a vector. The stack will be transferred through the RNN in the hidden state. We include the depth of each node in its positional embedding, which will help us determine which of the two operations (push or pop) to use.
@@
[…]
Finally, note that the memory rank of any tree with $n$ nodes is bounded by $\log(n)$ (see Lemma C.2). In fact, for a given tree structure $T$, the hidden dimension of the RNN needed to solve all CRQs with this structure is bounded by $O(d\cdot\textbf{mr}(T))$ using our construction. We next prove that the memory-rank also provides a lower bound on the required size of the hidden state of the RNN to solve CRQs:
@@
The proof is given in Section C.3. Combining Theorem 5.4 and Theorem 5.5, we see that the memory-rank provides a complete characterization of the memory required to solve CRQs using RNNs.
@@ 6 Solution Using Shallow Transformers with Chain of Thought
The proof idea is to solve all the sub-questions of the CRQ from the bottom-up. Each generated CoT token will be a solution to one of the sub-questions represented by a non-leaf node. Each node token contains two sets of positional encoding. The first one represents the tree structure and is similar to the positional encoding from Theorem 4.1. The second one encodes the position in the reverse BFS ordering of the nodes. The role of this embedding is so that at each operation of the transformer, the last token that was generated will attend to the next token in the reverse BFS order. This way, all the sub-question in the same layer are solved one after the other, before moving on to the next layer. The construction of the transformer itself contains $2$-layers of self-attention. The first layer is similar to a single layer of the construction from Theorem 4.1. Its role is to solve a single layer of sub-questions. The second layer make sure that the last generated token can only attend to the next token in the reverse BFS order, using the designated positional encodings.
[…]
@@
7 Generalization to Different Tasks
In our definition of CRQs (Definition 3.1) each sub-question is defined as the argmax over inner products. There are two main reasons for this choice: (1) The argmax function aligns well with the attention layers in transformers, allowing for simple constructions, and (2) CRQ is a flexible framework that includes natural problems like Boolean expressions as special cases. However, the CRQ framework can be modified to include many more tasks, which we term general CRQ (Definition F.1). Our upper bounds can be adapted to general CRQs. Namely, consider $k$ operators $f_{1},\dots,f_{k}:\mathbb{R}^{3d}\rightarrow\mathbb{R}^{d}$, where each operator can be implemented by a ReLU network. We can now define CRQs over binary trees, where each non-leaf $v$ node is labeled by both a vector $\mathbf{x}_{v}$ and an operator $f_{i}$ for some $i\in[k]$, while leaf nodes are labeled only by a vector. The answer to the sub-question of a node $v$ with leaf children $u_{1},u_{2}$ will be $f_{i}(\mathbf{x}_{v},\mathbf{x}_{u_{1}},\mathbf{x}_{u_{2}})$.
To give a concrete example of the generality of this formalism, it allows to represent arithmetic problems. Let $d=1$ and consider the operators $f_{+},f_{-},f_{\times},f_{\div}$ defined as $f_{+}(\mathbf{x}_{v},\mathbf{x}_{u_{1}},\mathbf{x}_{u_{2}})=\mathbf{x}_{u_{1}}+\mathbf{x}_{u_{2}}$ (ignoring the vector $\mathbf{x}_{v}$) and analogously for the other operators. Now, an arbitrary arithmetic formula can be parsed into a tree and reformulated as a general CRQ. In Appendix F we provide a more extensive discussion about these generalizations, and explain how our constructions can be readily adapted to solve them.
@@ 8 Discussion and Future Work
In this work we present a general framework of problems which we term compositional reasoning questions. We show the depth trade-offs of solving CRQs using deep transformers, RNNs, and shallow transformers with CoT. Our results indicate that although transformers are highly parallelizable, they must be deep being able to solve CRQs with depth scaling with the depth of the tree of CRQs, this depth requirement is necessary. On the other hand, RNNs and CoT can solve CRQs with constant depth, although their solution is not parallelizable, and requires sequential computation.
@@ Acknowledgments
This work was partially supported by the Alfred P. Sloan Foundation, awards NSF RI-1816753, NSF CAREER CIF 1845360, NSF CHS-1901091, NSF DMS-MoDL 2134216, and an NSF Graduate Research Fellowship. We thank Will Merrill for useful discussions while this work was being completed.
Appendix A Proofs from Section 4
A.1 Proof of Theorem 4.1
@@ Appendix A Proofs from Section 4
Theorem 4.1.
For any CRQ with up to $n$ nodes, there exist at most $|Y|\leq n$ possible queries. By Lemma A.1 there exist vectors $\mathbf{z}_{1},\dots,\mathbf{z}_{n}\in\{\pm 1\}^{4\log(n)}$ such that $|\left\langle\mathbf{z}_{i},\mathbf{z}_{j}\right\rangle|\leq 3\log(n)$ for any $i\neq j$.
@@
We construct the weights of the transformer so it will work as follows: In the first layer, each token representing an intermediate node will attend either to itself, if its depth is higher than $L-1$, or to its children, if its depth is exactly $L-1$. This means that only the tokens of nodes in depth $L-1$ will change while all other tokens of intermediate nodes will remain the same. Using the positional encodings, the tokens at depth $L-1$ will attend only to their children, not to themselves, and after the self-attention layer (before applying the $V$ matrix) will be equal to the token representing their child with the highest correlation to their query. We now use the $V$ matrix, residual connection and MLP so that the tokens with depth $L-1$ will have a similar embedding to the leaves tokens, and decrease the depth of all other intermediate tokens. We also make sure that the flag after the first layer will be equal to $1$ for nodes of depth $L-2$. We now apply a similar layer exactly $L$ times, so that in the final layer the output of the token representing the root will be equal to the solution to the CRQ.
@@
    If the entries in places $12d+4\log(n)+$ 1 to $2d+8\log(n)$ are zero, then $\mathbf{t}_{v}=\boldsymbol{0}$.
@@
    If the last entry of $\mathbf{t}_{v}$ is equal to $1$, then replace the entries in places $12d+$ 1 until $2d+4\log(n)$ (which used to hold $\mathbf{z}_{v}$) with the entries in places $12d+4\log(n)+$ 1 until $2d+8\log(n)$ (which used to hold $\mathbf{z}_{\mathcal{P}(v)}$). Then, replace the entries in places $12d+4\log(n)+$ 1 until $2d+8\log(n)$ with zeroes. This makes it a leaf embedding.
@@
    Add $1$ to the entry in place $12d+8\log(n)+$ 1. Then, if this entry is equal to $L-1$, replace the last entry with 1, if this entry is equal to $L$ replace the last two entries with zeroes.
@@
that can be implemented by the first layer of an MLP. We concatenate the output of this function on the $12d+4\log(n)+$ 1-th coordinate of the input, to the input itself. Then, we use the output layer to apply the function $\tilde{\mathbf{x}}\mapsto\boldsymbol{0}_{k}\cdot(\tilde{\mathbf{x}})_{k+1}+\tilde{\mathbf{x}}\cdot(1-(\tilde{\mathbf{x}})_{k+1})$, where $\tilde{\mathbf{x}}$ is the input concatenated with the additional coordinate of the output of $f$, and $k=2d+8\log(n)+2$. use similar constructions to implement the rest of the operations.
@@
[…]
@@
    If $\varphi$ is a Boolean formula, then ($(\neg\varphi)$ is a Boolean formula.
@@
The goal in BFEP, is given a Boolean formula to evaluate whether it is true or false (i.e. outputs $1$ or 00). We will construct a translation function $f$ from Boolean formulas to CRQ, and specifically to a tree structure where each node is given a specific value. We will use vectors in $\mathbb{R}^{2}$, while in fact all the possible values of the vectors in the CRQ will be in $\{0,\pm 1,\pm 2\}$. We define the vectors $\mathbf{t}_{1}=\begin{pmatrix}1\\ 0\end{pmatrix}$ and $\mathbf{t}_{0}=\begin{pmatrix}0\\ 1\end{pmatrix}$. The translation function is defined recursively as follows:
@@
    Given two Boolean formulas $\varphi_{1},\varphi_{2}$ we define $f((\varphi_{1}\land\varphi_{2}))$ and $f((\varphi_{1}\lor\varphi_{2}))$ as the tree in Figure 3 (right), where for $\land$ we use $\mathbf{t}_{0}$ for the root, and for $\lor$ we use $\mathbf{t}_{1}$.
@@
    Given a Boolean formulas $\varphi$ we define $f((\neg\varphi))$ as the tree in Figure 3 (left).
@@
[…]
$r_{i}=\sum_{j<i}\mathds{1}\big{[}\varphi_{j}=`(\text{'}\big{]}-\sum_{j<i}\mathds{1}\big{[}\varphi_{j}=`)\text{'}\big{]}~{}.$ ’ ] . (8)
[…]

### v2

@@ 4.2 Constant depth transformers cannot solve CRQs
We will now show that solving CRQs using a constant size transformer cannot be done, conditional on the assumption that ${\mathsf{TC}}^{0}\neq{\mathsf{NC}}^{1}$. Our main result in this subsection is a reduction from Binary Formula Evaluation Problem (BFEP), which is known to be ${\mathsf{NC}}^{1}$-complete (see Buss (1987)), to solving CRQ.
@@
The CRQ problem over a finite alphabet is ${\mathsf{NC}}^{1}$-hard.
For a full proof, see Section˜A.2. The reduction is straightforward, and resembles the reduction in Feng et al. (2024) from BFEP to arithmetic problems. The main idea is to define the vectors $\mathbf{t}_{0}=\mathbf{t}_{\land}=\begin{pmatrix}0\\ 1\end{pmatrix}$ and $\mathbf{t}_{1}=\mathbf{t}_{\lor}=\begin{pmatrix}1\\ 0\end{pmatrix}$ which correspond to 0 and $1$ in Boolean formulas. Now note that taking argmax of the dot product with $\mathbf{t}_{\land}$ corresponds to the $\land$ operation, and the argmax of the dot product with $\mathbf{t}_{\lor}$ corresponds to the $\lor$ operation. The operation of $\lnot$ is slightly more intricate and defined in Figure˜3. All the operations in the reduction can be done using ${\mathsf{TC}}^{0}$ circuits, which means that the CRQ problem is ${\mathsf{NC}}^{1}$-hard.
It was shown in Merrill and Sabharwal (2023c) that transformers with constant depth, polynomial size and logarithmic bit- precision (all w.r.t $n$) are in ${\mathsf{TC}}^{0}$. Thus, the following theorem follows immediately from Lemma˜4.2:
@@
Under the assumption that ${\mathsf{TC}}^{0}\neq{\mathsf{NC}}^{1}$, for any $L\in\mathbb{N}$ and polynomial $P(x)$, there exists $n\in\mathbb{N}$ such that no transformer with depth $L$, a number of parameters that is smaller than $P(n)$ and $O(\log(n))$ bit-precision can solve all CRQs of size $n$.
Combining the two results of this section, we see that transformers have the power to solve the CRQ problem for each layer efficiently, however the number of layers of the transformer depends on the size of the tree. This dependence is also mandatory under the assumption that ${\mathsf{TC}}^{0}\neq{\mathsf{NC}}^{1}$. One concrete example, which will be relevant in the next section, is a balanced binary tree with $n$ nodes, and $\log(n)$ depth. There exists a transformer that can solve all CRQs of this shape with $\log(n)$ layers, but not with a constant number of layers. In the next sections we will provide alternative approaches to solve this problem with different models using constant depth.
We note that arithmetic problems, as presented in Feng et al. (2024), can also be embedded as certain generalized CRQs. This embedding is presented in Appendix˜F. Thus, both Theorem˜4.1 and Theorem˜4.3 can be applied to arithmetic problems, if they are presented in a tree structure. It is in general possible to apply a preprocessing procedure using a depth $O(\log(n))$ transformer to turn an arithmetic problem presented in a sequential form, into a tree structure which captures the order of arithmetic operations. Note that the above theorem readily extends to transformers with layer normalization and attention masking, since it relies on a result for Merrill and Sabharwal (2023c) that includes both elements.
@@ 5 Solution Using Shallow RNNs
In this section we will show that under certain assumptions, RNNs can solve CRQs. The proofs for this section appear in Appendix˜C. We first define RNNs in the following way:
@@ 5.1 The order of the nodes matters
Our next result shows that there are bad orderings which force RNNs to have large hidden dimension:
@@
The proof is given in Section˜C.1. Note that Theorem˜5.2 doesn’t depend on the depth of the RNN. Note that it is easy to construct an RNN with a hidden dimension of size $O(n)$ that solves all CRQs. The idea is just to encode all the nodes in the hidden state, which can be done if we have enough memory, and then use a large enough neural network to solve the CRQ. Constructing such a neural network is possible by the universal approximation property (Cybenko, 1989; Leshno et al., 1993). Since the number of possible outputs is finite (because the alphabet is finite), we can approximate the solution up to a small enough accuracy and then threshold over the output.
The proof of Theorem˜5.2 uses a communication complexity argument. Specifically, we use a reduction from the set disjointness problem (see ˜C.1 for a formal definition). We then construct a communication protocol between two parties, where each one of them has knowledge of only half of the inputs. When using the bad ordering of the nodes, this communication protocol forces the first party to encode all of its inputs into the hidden state when passing it to the second party.
Feng et al. (2024) proved that a specific RNNs construction requires a hidden state of size $\Omega(n)$ to solve arithmetic problems. Our results are stronger in that: (1) Our lower bound applies to any possible construction, so long as the bad order is used, and (2) We next show that the $\Omega(n)$ memory requirement can be alleviated by re-ordering the inputs.
@@ 5.2 Memory-rank sort
In this subsection, we restrict ourselves to CRQs with full binary trees for simplicity; that is, the degree of each node is 0 or $2$. In Appendix˜D, we extend the results from this section to non-binary trees, but note that every CRQ can be converted into an equivalent CRQ with a binary tree and at most twice as many notes (see Section˜D.1).
@@
Intuitively, the memory rank is the smallest possible stack size needed for a stack machine to solve the CRQ if we are allowed to pick the order of the nodes. A stack machine is an automaton augmented with a stack that it can uses as memory. It processes the inputs sequentially, optionally pushing and popping a finite number of elements from its stack at each step and performing computations with the resulting vectors. We should order the inputs according to a post-ordering depth-first search of the CRQ, meaning that the position of a node in the ordering corresponds to when the depth first search last visits it. This ordering akin to reverse Polish notation, which notates an arithmetic expression like (2+3)6(2+3)*6 as the following sequence: 2 3+ 62\,3\,+\,6\,*. By ordering the inputs this way, we ensure that whenever the automaton reads a non-leaf node, the top two elements on the stack are the solutions to the CRQs defined by that node’s two children. It can simply pop them off, compute the solution to the current node, and push the result back onto the stack to save it for later.
Even using a depth-first search ordering, the size of the stack used by this machine could grow as large as $n$ in the worst case. However, by ordering the children of each node carefully, we can reduce the stack depth to only $\log n$. To compute the CRQ corresponding to a given node $v$, we (1) solve for one of its children, (2) save this result on the stack, then (3) solve for the second child. If we solve $c_{1}(v)$ first, then the largest stack depth during the first stage is $\textbf{mr}(c_{1}(v))$ and the largest stack depth during the second stage is $\textbf{mr}(c_{2}(v))+1$. The overall largest stack depth is the maximum of these two. Thus, to use as little memory as possible, we should always start with the child whose memory rank is larger. This sorting algorithm is defined formally in Algorithm˜1. See Figure˜2 for an example of memory-rank sort of a tree. In Theorem˜5.4, we will simulate this stack machine using a RNN with hidden dimension of size $\textbf{mr}(T)$.
@@
Figure 2: Memory-rank sorting of a tree. Each node is labeled with its memory rank, which ranges from 0 to 2 in this case. The number below each node is its ordering according to Algorithm˜1. Note that we first traverse the right branch of the tree, since the right child of the root has a higher memory-rank than the left child.
@@ 5.3 Shallow RNNs can solve CRQs
For any $n\in\mathbb{N}$ there exists an RNN with $5$ layers and hidden dimension $O(d\log(n))$ that solves any CRQ with a binary tree, if the nodes are ordered by Algorithm˜1
The proof is given in Section˜C.2. The main idea is to simulate a stack machine with the RNN. The stack will contain the vectors $\mathbf{x}_{i}$ that are needed for future calculations, and will have two possible operations: (1) pop out vectors, and (2) push a vector. The stack will be transferred through the RNN in the hidden state. We include the depth of each node in its positional embedding, which will help us determine which of the two operations (push or pop) to use.
@@
Note that in Theorem˜4.1, we also needed the embedding dimension of the transformer to be $\Omega(\log(n))$, but for a very different reason. There, the transformer is invariant to re-ordering of the input tokens, and since the output does depend on the ordering (specifically, the structure of the tree) we would need $\Omega(\log(n))$ bits just to label the tokens from $1$ to $n$. For RNNs, the inputs are already ordered, and we even defined a specific ordering algorithm that aligns with our task. The reason that in Theorem˜5.4 we need the size of the hidden dimension to be $\Omega(\log(n))$ is because in order to solve CRQs, we need to store some of the inner calculations (i.e. answers to the sub-questions) while we solve others. We emphasize that this hardness is not captured by the fact that solving CRQs is ${\mathsf{NC}}^{1}$-hard. There are other ${\mathsf{NC}}^{1}$-hard problems that can be solved with a constant size RNN, such as the word problem on $S_{5}$ (see Definition 3.1 and Theorem 5.1 in Merrill et al. (2024)).
Finally, note that the memory rank of any tree with $n$ nodes is bounded by $\log(n)$ (see Lemma˜C.2). In fact, for a given tree structure $T$, the hidden dimension of the RNN needed to solve all CRQs with this structure is bounded by $O(d\cdot\textbf{mr}(T))$ using our construction. We next prove that the memory-rank also provides a lower bound on the required size of the hidden dimension:
@@
The proof is given in Section˜C.3. Combining Theorem˜5.4 and Theorem˜5.5, we see that the memory-rank gives a complete characterization of the memory needed to solve CRQs using RNNs.
@@ 6 Solution Using Shallow Transformers with Chain of Thought
The proof idea is to solve all the sub-questions of the CRQ from the bottom-up. Each generated CoT token will be a solution to one of the sub-questions represented by a non-leaf node. Each node token contains two sets of positional encoding. The first one represents the tree structure and is similar to the positional encoding from Theorem˜4.1. The second one encodes the position in the reverse BFS ordering of the nodes. The role of this embedding is so that at each operation of the transformer, the last token that was generated will attend to the next token in the reverse BFS order. This way, all the sub-question in the same layer are solved one after the other, before moving on to the next layer. The construction of the transformer itself contains $2$-layers of self-attention. The first layer is similar to a single layer of the construction from Theorem˜4.1. Its role is to solve a single layer of sub-questions. The second layer make sure that the last generated token can only attend to the next token in the reverse BFS order, using the designated positional encodings.
Our work leaves open the question of whether a constant-depth transformer can solve CRQs by generating $o(n)$ CoT tokens. In Li et al. (2024) it was shown that transformers with constant depth that generate $O(\log(n))$ CoT tokens lie in ${\mathsf{TC}}^{0}$ (see also Theorem 4 of Merrill and Sabharwal (2025)). By Lemma˜4.2 we know that solving CRQs is ${\mathsf{NC}}^{1}$-hard. Thus, under the assumption that ${\mathsf{TC}}^{0}\neq{\mathsf{NC}}^{1}$, we know that transformers need more than logarithmically many CoT tokens to solve CRQs. However, the gap between $\Omega(\log(n))$ and our solution with $O(n)$ CoT tokens is still open. We also note that by Amiri et al. (2025), in the finite precision case, a lower bound of $\Omega(n)$ CoT tokens for solving CRQs can be derived. Hence, in this case, our construction is asymptotically optimal. In the log-precision case (which is the main setting of this paper), the question of whether our construction is optimal remains open.
@@ 7 Discussion and Future Work
In this paper, we introduce the framework of compositional reasoning questions. We describe the trade-offs inherent in solving CRQs with deep transformers, RNNs, and shallow transformers with CoT. Our results indicate that although transformers are highly parallelizable, they must be deep to be able to solve CRQs; in particular, the depth must scale with the depth of the CRQ tree. In contrast, RNNs and CoT can solve CRQs with constant depth, but their operation is not parallelizable. Transformers require quadratic computational cost, but RNNs can solve CRQs in nearly linear time. Finally, transformers with CoT solve CRQs using a logarithmic number of parameters, while the other models require the number of parameters to be linear in $n$ in the worst case.
Our results have a broader significance for the grand challenge of improving LLM reasoning. The three architectures we consider correspond to three schools of thought about how to make progress on this problem, each of which has received support from the theoretical literature:
  1. 1.
    Scale up transformers by, e.g., adding more layers, and greater capabilities will continue to emerge. Transformers with depth $\log(n)$ can do k-hop induction and graph algorithms (Merrill and Sabharwal, 2025; Sanford et al., 2024b, c).
  2. 2.
    Devise better architectures, such as state space RNNs or hybrid models, because attention is not all you need to learn and reason efficiently. RNNs can more easily do state tracking tasks like recognizing regular languages (Merrill et al., 2024).
  3. 3.
    Introduce test-time scaling via chain of thought, opening a new axis of the design space. This enables dynamic programming and simulating Turing machines (Feng et al., 2024; Merrill and Sabharwal, 2023b).
In previous work, the school that looks most promising depends on which model problem they chose to study. CRQs are just as fundamental as any of the other problems studied in the literature. However, when using CRQs as a theoretical yardstick of reasoning, we find that none of the schools is a clear winner and none is a clear loser. Thus, the theory of representational capacity does not give a single answer to the question of how to improve LLM reasoning. Perhaps some blend of the three schools or a fourth school will emerge that dominates the others, but for now the answer is equivocal and contingent, even for a single task.
@@ Acknowledgments
NA was supported by NSF award 2234660. We thank Will Merrill for helpful discussions related to this work and for drawing our attention to the Boolean formula evaluation problem.
Appendix A Proofs from Section˜4
A.1 Proof of Theorem˜4.1
@@ Appendix A Proofs from Section˜4
Theorem˜4.1.
For any CRQ with up to $n$ nodes, there exist at most $|Y|\leq n$ possible queries. By Lemma˜A.1 there exist vectors $\mathbf{z}_{1},\dots,\mathbf{z}_{n}\in\{\pm 1\}^{4\log(n)}$ such that $|\left\langle\mathbf{z}_{i},\mathbf{z}_{j}\right\rangle|\leq 3\log(n)$ for any $i\neq j$.
@@
We construct the weights of the transformer so it will work as follows: In the first layer, each token representing an intermediate node will attend either to itself, if its depth is higher than L1L-1, or to its children, if its depth is exactly L1L-1. This means that only the tokens of nodes in depth L1L-1 will change while all other tokens of intermediate nodes will remain the same. Using the positional encodings, the tokens at depth L1L-1 will attend only to their children, not to themselves, and after the self-attention layer (before applying the $V$ matrix) will be equal to the token representing their child with the highest correlation to their query. We now use the $V$ matrix, residual connection and MLP so that the tokens with depth L1L-1 will have a similar embedding to the leaves tokens, and decrease the depth of all other intermediate tokens. We also make sure that the flag after the first layer will be equal to $1$ for nodes of depth L2L-2. We now apply a similar layer exactly $L$ times, so that in the final layer the output of the token representing the root will be equal to the solution to the CRQ.
@@
    If the entries in places $2d+4\log(n)+1$ to $2d+8\log(n)$ are zero, then $\mathbf{t}_{v}=\boldsymbol{0}$.
@@
    If the last entry of $\mathbf{t}_{v}$ is equal to $1$, then replace the entries in places $2d+1$ until $2d+4\log(n)$ (which used to hold $\mathbf{z}_{v}$) with the entries in places $2d+4\log(n)+1$ until $2d+8\log(n)$ (which used to hold $\mathbf{z}_{\mathcal{P}(v)}$). Then, replace the entries in places $2d+4\log(n)+1$ until $2d+8\log(n)$ with zeroes. This makes it a leaf embedding.
@@
    Add $1$ to the entry in place $2d+8\log(n)+1$. Then, if this entry is equal to L1L-1, replace the last entry with 1, if this entry is equal to $L$ replace the last two entries with zeroes.
@@
that can be implemented by the first layer of an MLP. We concatenate the output of this function on the $2d+4\log(n)+1$-th coordinate of the input, to the input itself. Then, we use the output layer to apply the function $\tilde{\mathbf{x}}\mapsto\boldsymbol{0}_{k}\cdot(\tilde{\mathbf{x}})_{k+1}+\tilde{\mathbf{x}}\cdot(1-(\tilde{\mathbf{x}})_{k+1})$, where $\tilde{\mathbf{x}}$ is the input concatenated with the additional coordinate of the output of $f$, and $k=2d+8\log(n)+2$. use similar constructions to implement the rest of the operations.
@@
Our proof relies on a reduction to the Boolean Formula Evaluation Problem (BFEP). It was shown in Buss [1987], that this problem is ${\mathsf{NC}}^{1}$-complete, thus a reduction of BFEP to CRQ would finish the proof. The BFEP is defined recursively over the alphabet: $\Sigma=\{1,0,\lor,\land,\neg,(,)\}$ in the following way:
@@
    If $\varphi$ is a Boolean formula, then $(\neg\varphi)$ is a Boolean formula.
@@
The goal in BFEP, is given a Boolean formula to evaluate whether it is true or false (i.e. outputs $1$ or 0). We will construct a translation function $f$ from Boolean formulas to CRQ, and specifically to a tree structure where each node is given a specific value. We will use vectors in $\mathbb{R}^{2}$, while in fact all the possible values of the vectors in the CRQ will be in $\{0,\pm 1,\pm 2\}$. We define the vectors $\mathbf{t}_{1}=\begin{pmatrix}1\\ 0\end{pmatrix}$ and $\mathbf{t}_{0}=\begin{pmatrix}0\\ 1\end{pmatrix}$. The translation function is defined recursively as follows:
@@
    Given two Boolean formulas $\varphi_{1},\varphi_{2}$ we define $f((\varphi_{1}\land\varphi_{2}))$ and $f((\varphi_{1}\lor\varphi_{2}))$ as the tree in Figure˜3 (right), where for $\land$ we use $\mathbf{t}_{0}$ for the root, and for $\lor$ we use $\mathbf{t}_{1}$.
@@
    Given a Boolean formulas $\varphi$ we define $f((\neg\varphi))$ as the tree in Figure˜3 (left).
We will first show that this construction can be done using ${\mathsf{TC}}^{0}$ circuits, and then show its correctness. Given a Boolean formula of length $s$, it is clear that the construction above creates a tree with at most $O(|s|)$ nodes, where the values of each nodes is a $2$-d vectors with entries in $\{0,\pm 1,\pm 2\}$. Thus, the total number of bits requires for the construction is also bounded by $O(|s|)$.
The translation works recursively over the logical operators $\neg,\land,\lor$, and to make the construction work we need to know in which order to translate them. Namely, the root should begin with the outmost logical operator, its children the second to outmost operators etc., until we reach either 0 or $1$, which should be translated to leaves in the tree. We will show that this order can be determined using ${\mathsf{TC}}^{0}$ circuits. Given a Boolean formula $\varphi$ we define by $\varphi_{i}$ its character in place $i$. We define:
@@
It is easy to see that $r_{i}$ can be calculated using ${\mathsf{TC}}^{0}$ circuits. Now, the order in which the recursion works is in increasing order w.r.t $r_{i}$ over the logical operators. Namely, begin with the logical operator in place $i$ where $r_{i}=0$, continue with $r_{i}=\pm 2$ an so forth.
For the correctness of the reduction, it is an easy calculation to see that given $t_{i}$ for $i\in\{0,1\}$ the construction for $\neg\varphi$ outputs $t_{1-i}$ with the definition of a CRQ. Also, for $\land$ and $\lor$ the construction outputs correctly $t_{i}\land t_{j}$ and $t_{i}\lor t_{j}$ for $i,j\in\{0,1\}$. Therfore, this construction forms a reduction from the BFEP to CRQ, and thus CRQ is ${\mathsf{NC}}^{1}$-hard.